solve by either the substitution or the addition-orsubtraction method: 2m=n-3 and 3m=2n+9
Answer by ewatrrr(24785)

 
Hi
solving by substitution
2m = n-3 OR n = 2m+3
3m = 2n+9
3m = 2(2m+3) + 9
3m = 4m+6+9
-15 = m and n = -27 ( n = 2*-15 + 3)